A bookstore has a clearance shelf with $3 books and $5 books. On Sat the bookstore sold 96 of the clearance books for a total of $400. How many of the $3 books were sold?
Found 2 solutions by ikleyn, jorel555:
Answer by ikleyn(52787)

.
A bookstore has a clearance shelf with $3 books and $5 books. On Sat the bookstore sold 96 of the clearance books for a total of $400. How many of the $3 books were sold?
~~~~~~~~~~~~~~~~~~~~~~~~~~~

 x +  y =  96,    (1)   (the total amount of books)
3x + 5y = 400.    (2)   (the "value" equation).

Solve by any method you know.


Answer by jorel555(1290)

Let n be the amount of $5 books sold. Then:
5n+3(96-n)=400
2n+288=400
2n=112
n=56
56 of the $5 books were sold; and 40 of the $3 books. ☺☺☺☺